what is (x^2+9x)(x^2-9x) multiplied
Answer by nyc_function(2741)

(x^2+9x)(x^2-9x) =

USE THE FOIL METHOD.

x^2 * x^2 = x^4

x^2 * - 9x = -9x^3

9x * x^2 = 9x^3

9x * -9x = -81x^2

Put it all together and combine like terms.

x^4 - 9x^3 + 9x^3 - 81x^2

x^4 - 81x^2

We now factor.

x^2(x^2 - 81)

We now factor what lies in the parentheses.

Final Answer: x^2 (x - 9) (x + 9)